OpenStudy (anonymous):

Robert purchased $2,000 worth of U.S. savings bonds. If bonds are sold in $50 or $100 denominations only, which of the following CANNOT be the number of U.S. savings bonds that Robert purchased? (1 point) 29 27 30 40 50

OpenStudy (anonymous):

let's say a = no. fo 50$ bonds and b = no. fo 100$ bonds 50a+100b=2000 => a+2b=40 then you have: a) a+b = 29 => b=11, a = 18 do the same for b, c, d, e The number of bonds cannot be 50 a+b=50 a+2b=40 => b=-10, a=60 ( b it's a negative number)

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

50 lol